OpenStudy (anonymous):

How would I write an equation of quadratic equation with zeros of (–2, 0) and (5, 0), and passes through the point: (7, –4)

OpenStudy (jamesj):

Nice question. Well, remember (or not :-) ) that a quadratic equation with roots a and b can always be factorized in the form (x-a)(x-b) because if (x-a)(x-b) = 0, then x = a or x = b But notice this isn't the only quadratic that has roots a and b. So does 2(x-a)(x-b) and -5(x-a)(x-b) Hence in general for any c not equal to zero the quadratic c(x-a)(x-b) has roots at x = a or b. That's the theory. Now for your problem you are given these two points on the graph: (-2,0) and (5,0) and told they are zeros, i.e., they are roots. So the GENERAL form of the quadratic with those roots is y = c(x+2)(x-5) --- ** Now you need find the value for c which makes the graph of this quadratic pass through (7,-4). Hence substitute into the equation ** above x = 7 and y = -4 and solve for c.

OpenStudy (anonymous):

So like this: -4=c(7+2)(7-5) and then solve for c?

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

I got y=(-2/9)(x+2)(x-5) as the answer

OpenStudy (jamesj):

I'd expand it out, as that's the usual form, but I think that looks right.